Munoz Chairs, Inc. makes two types of chairs. Model Diamond is a high-end product designed for professional offices. Model Gold is an economical product designed for family use. Jane Silva, the president, is worried about cut-throat price competition in the chairs market. Her company suffered a loss last quarter, an unprecedented event in its history. The company’s accountant prepared the following cost data for Ms. Silva:

Direct Cost per Unit Model Diamond (D) Model Gold (G)
Direct materials $ 21.10 per unit $ 10.60 per unit
Direct labor $ 18.60 /hour × 2.00 hours production time $ 18.60 /hour × 1.00 hour production time
Category Estimated Cost Cost Driver Use of Cost Driver
Unit level $ 379,500 Number of units D: 15,000 units; G: 40,000 units
Batch level 1,225,000 Number of setups D: 110 setups; G: 240 setups
Product level 558,000 Number of TV commercials D: 6; G: 12
Facility level 676,000 Number of machine hours D: 2,000 hours; G: 3,200 hours
Total $ 2,838,500


The market price for office chairs comparable to Model Diamond is $121 and to Model Gold is $74.

Required

  1. Compute the cost per unit for both products.

  2. Dan Barker, the chief engineer, told Ms. Silva that the company is currently making 136 units of Model Diamond per batch and 167 units of Model Gold per batch. He suggests doubling the batch sizes to cut the number of setups in half, thereby reducing the setup cost by 50 percent. Compute the cost per unit for each product if Ms. Silva adopts his suggestion.
Answer #1

a) Calculation of cost per unit :

Product Model Diamond :

Note: Total number of unit of product model diamond produced is 15,000 units

Particulars

Amount (cost per unit) $

Direct Material

21.10

Direct labour ( Number of hours * per hour rate)

=( 2 * 18.60)

37.2

Direct Overhead

Unit level ( Total unit level cost * nos. of units of model diamond / Total nos. Of unit produced)

=( 379,500 * 15,000 / 55,000) = $103,500

Therefore unit level cost per unit =( Total cost / nos. Of unit of model diamond produced)

=( $103,500 / 15000) =

6.9

Batch level (Total batch level cost * nos. Of setup of model diamond / Total nos. Of setup)

= (1,225,000 * 110 / 350) = $385,000

Therefore batch level cost per unit =( Total cost / nos. Of unit of model diamond produced)

= ($385,000 / 15000) =

25.67

Product level (Total production level cost * nos. Of TV commercial of diamond / total nos. Of TV commercials)

= (558,000 * 6 / 18) = $186,000

Therefore product level cost per unit =( Total cost / nos. Of unit of model diamond produced)

=( $186,000 / 15,000) =

12.4

Facility level (Total facility level cost * nos. Of machine hours of diamond / total nos of machine hour)

= (676,000 * 2000 / 5200) = $260,000

Therefore facility level cost per unit =( Total cost / nos. Of unit of model diamond produced)

= ($260,000 / 15,000) =

17.33

Total cost per unit

120.60

Product Model Gold :

Note: Total number of unit of product model diamond produced is 40,000 units

Particulars

Amount (cost per unit) $

Direct Material

10.60

Direct labour ( Number of hours * per hour rate)

=( 1 * 18.60)

18.60

Direct Overhead

Unit level ( Total unit level cost * nos. of units of model gold / Total nos. Of unit produced)

=( 379,500 * 40,000 / 55,000) = $276,000

Therefore unit level cost per unit =( Total cost / nos. Of unit of model gold produced)

=( $276,000 / 40000) =

6.9

Batch level (Total batch level cost * nos. Of setup of model gold / Total nos. Of setup)

= (1,225,000 * 240 / 350) = $840,000

Therefore batch level cost per unit =( Total cost / nos. Of unit of model gold produced)

= ($840,000 / 40,000) =

21

Product level (Total production level cost * nos. Of TV commercial of gold / total nos. Of TV commercials)

= (558,000 * 12 / 18) = $372,000

Therefore product level cost per unit =( Total cost / nos. Of unit of model gold produced)

=( $372,000 / 40,000) =

9.3

Facility level (Total facility level cost * nos. Of machine hours of gold / total nos of machine hour)

= (676,000 * 3200 / 5200) = $416,000

Therefore facility level cost per unit =( Total cost / nos. Of unit of model gold produced)

= ($416,000 / 40,000) =

10.4

Total cost per unit

76.80

b) If Ms. Silva adopts the suggestion of Dan Barker , the chief enginee rand doubles the batch size the cost per unit of both the product model will be as:-

Revised total cost will be :

Total Batch cost = $1,225,000 * 50% = $612500

Cost driver = nos. Of setup

Use of cost driver :

For Diamond = 110 setups * 50% = 55 setup

For gold = 240 setups * 50% = 120 setups

Therefore, revised Batch cost

For Diamond = (Total batch level cost * nos. Of setup of model diamond / Total nos. Of setups)

= ($612,500 * 55 / 175) = $192,500

Therefore batch level cost per unit =( Total cost / nos. Of unit of model diamond produced)

= ($192,500 / 15000) =$12.83 per unit.

For Gold = (Total batch level cost * nos. Of setup of model gold / Total nos. Of setup)

= ($612,500 * 120 / 175) = $420,000

Therefore batch level cost per unit =( Total cost / nos. Of unit of model gold produced)

= ($420,000 / 40,000) = $10.5 per unit

Remainning cost will not change.

Therefore Revised cost per unit of diamond = $120.60 - $25.37 + $12.83 = $108.06 per unit.

Therefore Revised cost per unit of gold = $76.80 - $21 + $10.5 = $66.30 per unit.